### v1

# Quantum many-body analysis of a spin-2 bosons with two-body inelastic decay

Abstract
Bose-Einstein condensates (BECs) of $^{87}\textrm{Rb}$ atoms with a hyperfine spin of 2 are open quantum systems, where the atoms are lost through two-body inelastic collisions. In this dissipation process, a collision channel with total spin of 4 is forbidden by angular momentum conservation, which results in magnetization of the atoms remaining in the condensate. Here, we investigate the quantum many-body properties of spin-2 bosons that undergo two-body atomic loss. We show that the system finally reaches a steady state, which is a mixture of the states with maximum total spins. In addition, we find that a non-classical steady state can be obtained by quenching the quadratic Zeeman coefficient.
I Introduction
Atomic Bose-Einstein condensates (BECs) with spin degrees of freedom [1] have been produced experimentally by far-off-resonant optical traps [2]. The spin degrees of freedom enable a rich variety of ground-state phases [3, 4, 5] and topological excitations [6, 7]. For example, skyrmions, monopoles, and knots have been experimentally demonstrated [8, 9, 10, 11, 12].
Spin-2 BECs have attracted much interest because they exhibit more diverse magnetic phases than spin-1 BECs [13, 14, 15]. However, for the atomic species used in experiments (e.g., $^{87}\textrm{Rb}$), the spin-2 hyperfine state has a larger energy than the spin-1 state and is unstable against transition to the spin-1 state through binary collisions [16, 17, 18]. In such inelastic collisions, the hyperfine energy (e.g., $\simeq 6.8$ GHz for $^{87}\textrm{Rb}$) transforms to kinetic energy and the colliding atoms escape from the trap. Such a system with particle dissipation is an example of an open quantum system [19, 20, 21].
In the experiment reported in Ref. [21], magnetization was observed to dynamically develop in a spin-2 BEC of $^{87}\textrm{Rb}$ atoms, even though the spin-dependent contact interaction was not ferromagnetic. The mechanism involved in such magnetization is spin-dependent atomic loss, in which two-body inelastic collisions of particles with a total spin of 4 (=2+2=2+2) is prohibited because of angular momentum conservation (i.e., only atomic pairs having different spin directions are lost). Consequently, the remaining atoms tend to become magnetized in the same direction.
In Ref. [21], the experimental results have been confirmed by numerical calculations using the mean-field approximation. However, the mean-field picture is inappropriate for systems with a small number of particles. In fact, when the trap confinement is tight and all atoms occupy the same spatial wave function, the true ground state of a spin-1 BEC with antiferromagnetic interaction is a symmetry-preserving state referred to as a fragmented state [24, 22, 23]. Recently, such a quantum many-body state was experimentally observed in a BEC with $\sim 100$ atoms [25, 26].
In the present study, we perform quantum many-body analysis of spin-2 bosons that undergo two-body inelastic loss. We prepare an initial state in which all atoms occupy the $m=0$ magnetic sublevel, as in the experiment [21] and investigate the time evolution using the Lindblad master equation [27]. We show that the system eventually reaches a steady state, which is a statistical mixture of the states with maximum total spins (i.e., the final state is fully magnetized). We perform numerical simulations of the time evolution of the density matrix and investigate physical quantities in the final state. In the presence of the quadratic Zeeman effect, the transverse magnetization is found to be enhanced during the early stage of evolution of the system. When the quadratic Zeeman coefficient is quenched, the final steady state acquires a large transverse magnetization. This state exhibits a nonclassical magnetization distribution, indicating a Schrödinger-cat-like state.
The remainder of this paper is organized as follows. First, we formulate the open quantum system in Sec. II and describe the dynamics of the quantum state using the Lindblad master equation. In Sec. III, we focus on the steady states of this equation and discuss their general properties. In Sec. IV, we present the results of a numerical analysis of the dynamics of the system. The conclusions to this study are provided in Sec. V.
II Formulation of problem
II.1 Master equation with atomic loss
We consider spin-2 bosonic atoms confined in a trap potential $V(\bm{r})$. The field operators $\hat{\psi}_{m}(\bm{r})$ annihilate atoms in the magnetic sublevels m=2,1,0,1,2m=-2,-1,0,1,2, which satisfy the bosonic commutation relations $\hat{\psi}_{m}(\bm{r}),\hat{\psi}_{m^{\prime}}^{\dagger}(\bm{r}^{\prime})]=\delta(\bm{r}-\bm{r}^{\prime})\delta_{m,m^{\prime}}$. The non-intaracting component of the Hamiltonian is given by
$\hat{H}_{0}=\int\sum_{m=-2}^{2}\left\{\hat{\psi}^{\dagger}_{m}(\bm{r})\left[-\frac{\hbar^{2}\nabla^{2}}{2M}+V(\bm{r})\right]\hat{\psi}_{m}(\bm{r})\right$. (1)
$+qm^{2}\hat{\psi}^{\dagger}_{m}(\bm{r})\hat{\psi}_{m}(\bm{r})\bigg\}d\bm{r},$
where $M$ is the mass of an atom and $q$ is the quadratic Zeeman coefficient. Here, we ignore the linear Zeeman effect because it can be eliminated by the unitary transformation.
The interaction term $\hat{H}_{\textrm{int}}$ that represents the elastic collisions is written as
$\hat{H}_{\textrm{int}}=\int\sum_{\mathcal{F}=0,2,4}\frac{g_{\mathcal{F}}}{2}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})d\bm{r},$ (2)
where $\mathcal{F}$ is the total spin of the colliding atoms and $g_{\mathcal{F}}=4\pi\hbar^{2}a_{\mathcal{F}}/M$ with $a_{\mathcal{F}}$ being the $s$-wave scattering length of the collision channel with $\mathcal{F}$. The operator $\mathcal{\hat{A}}_{\mathcal{FM}}$ is defined as
$\hat{A}_{\mathcal{FM}}(\bm{r})=\sum_{m,m^{\prime}=-2}^{2}C_{m,m^{\prime}}^{\mathcal{FM}}\hat{\psi}_{m}(\bm{r})\hat{\psi}_{m^{\prime}}(\bm{r}),$ (3)
where $C_{m,m^{\prime}}^{\mathcal{FM}}$ is the Clebsch-Gordan coefficient. The Hamiltonian for the total system is given by $\hat{H}=\hat{H}_{0}+\hat{H}_{\textrm{int}}$.
Next, we consider two-body inelastic collisions, in which a transition from a hyperfine spin-2 state to a hyperfine spin-1 state occurs. The energy difference between hyperfine spin-2 and spin-1 states for $^{87}\textrm{Rb}$ is $\simeq 6.8$ GHz, and this energy is transferred to the kinetic energy of the colliding atoms, causing them to escape from the trap. The dynamics of such an open quantum system with atomic loss can be described by the density matrix $\hat{\rho}$ as
$\frac{\partial\hat{\rho}}{\partial t}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}]+\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}},$ (4)
where $(\partial\hat{\rho}/\partial t)_{\textrm{loss}}$ represents the atomic loss due to inelastic collisions. Assuming that the atoms escaping from the trap do not interact with other atoms, the atomic loss is described by a Markov process. In this case, $(\partial\hat{\rho}/\partial t)_{\textrm{loss}}$ in Eq. (4) is given in the Lindblad form as [20]
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}}=\int\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left[2\hat{A}_{\mathcal{FM}}(\bm{r})\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\right$. (5)
$\left.-\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})\hat{\rho}-\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})\right]d\bm{r},$
where $b_{\mathcal{F}}$ is the loss coefficient for a collision channel with total spin $\mathcal{F}$. Notably, for $\mathcal{F}=4$, inelastic collisions are forbidden (i.e., $b_{4}=0$) because of the conservation of spin angular momentum. The first term in the square bracket in Eq. (5) is called a jump term, which eliminates two particles from the system.
II.2 Previous results
Here, we briefly review the results of a previous study on a spin-2 $^{87}{\rm Rb}$ BEC with particle dissipation [21]. In this study, all atoms were initially in the $m=0$ state and the subsequent dynamics was investigated. It was observed that the system eventually became fully magnetized in the transverse direction despite the ferromagnetic state being energetically unfavorable [15]. This is due to the spin-dependent inelastic collisions ($b_{4}=0$); that is, atoms with different spin directions are selectively lost.
A numerical analysis was also performed in Ref. [21], in which Eq. (4) was reduced to the dissipative Gross-Pitaevskii equation. In the mean-field approximation, the macroscopic wave function $\psi_{m}(\bm{r})\equiv\textrm{Tr}[\hat{\psi}_{m}(\bm{r})\hat{\rho}$] obeys [20]
$i\hbar\frac{\partial\psi_{m}}{\partial t}=\left[-\frac{\hbar^{2}\nabla^{2}}{2M}+V(\bm{r})+qm^{2}\right]\psi_{m}(\bm{r})$ (6)
$+\sum_{\begin{subarray}{c}\mathcal{F}=0,2,4\\ \mathcal{M}=-\mathcal{F},...,\mathcal{F}\end{subarray}}\tilde{g}_{\mathcal{F}}\sum_{\begin{subarray}{c}i,j,k=-2\end{subarray}}^{2}C_{m,i}^{\mathcal{FM}}C_{j,k}^{\mathcal{FM}}\psi_{i}^{\star}(\bm{r})\psi_{j}(\bm{r})\psi_{k}(\bm{r}),$
where $\tilde{g}_{\mathcal{F}}=g_{\mathcal{F}}-i\hbar b_{\mathcal{F}}/2$ for $\mathcal{F}=0,2$, and $\tilde{g}_{4}=g_{4}$. In Ref. [21], the numerical solutions of Eq. (6) were found to agree well with the experimental results. However, the mean-field approximation is only valid for a large number of atoms and we do not use Eq. (6) in the present study.
II.3 Single-mode approximation
We consider a situation in which the trap is tight and the size of the system is much smaller than the spin healing length. In this case, all particles occupy the same spatial wave function with a uniform spin state and we can use the single-mode approximation. The field operators can be rewritten as
$\hat{\psi}_{m}(\bm{r})\simeq\hat{a}_{m}\psi_{\textrm{SMA}}(\bm{r}),$ (7)
where $\hat{a}_{m}$ are bosonic operators and the wave function $\psi_{\textrm{SMA}}(\bm{r})$ is the ground state of $\hbar^{2}\nabla^{2}/(2M)+V]\psi_{\rm SMA}=\varepsilon_{0}\psi_{\rm SMA}$ satisfying the normalization condition $\int|\psi_{\textrm{SMA}}(\bm{r})|^{2}d\bm{r}=1$. Substituting Eq. (7) into Eqs. (1) and (2), we obtain
$\hat{H}$= $\varepsilon_{0}\hat{N}+q\sum_{m=-2}^{2}m^{2}\hat{a}^{\dagger}_{m}\hat{a}_{m}$ (8)
$+\sum_{\mathcal{F}=0,2,4}\frac{g_{\mathcal{F}}}{2V^{\textrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}\hat{A}_{\mathcal{FM}}$,
where $\hat{N}=\sum_{m}\hat{a}_{m}^{\dagger}\hat{a}_{m}$, $V^{\rm eff}=\left[\int|\psi_{\textrm{SMA}}(\bm{r}))|^{4}d\bm{r}\right]^{-1}$, and
$\hat{A}_{\mathcal{FM}}=\sum_{m,m^{\prime}=-2}^{2}C_{m,m^{\prime}}^{\mathcal{FM}}\hat{a}_{m}\hat{a}_{m^{\prime}}.$ (9)
Similarly, substituting Eq. (7) into Eq. (5) gives
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}}=\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4V^{\textrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left(2\hat{A}_{\mathcal{FM}}\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}\right$. (10)
[…]
III GENERAL FORMS OF STEADY STATES
This section provides a general discussion of the steady state for the Lindblad equation (4) with a single-mode approximation. The steady states satisfy $\partial\hat{\rho}/\partial t=0$. In the present case, the unitary and dissipation terms on the right-hand side of Eq. (4) must each be zero in a steady state.
First, let us consider the condition for
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{unitary}}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}]=0.$ (11)
Denoting the energy eigenstates of $\hat{H}$ by $|n\rangle$ and their corresponding energy eigenvalues by $E_{n}$, we can generally write the density matrix as
$\hat{\rho}=\sum_{n,m}c_{nm}|n\rangle\langle m|,$ (12)
where $c_{nm}$ is a positive semidefinite Hermitian matrix with unit trace. Equation (11) can be written as
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{unitary}}$ $=\frac{1}{i\hbar}\sum_{n,m}(E_{n}-E_{m})c_{nm}|n\rangle\langle m|$ (13)
$=0.$
Therefore, the density matrix that satisfies Eq. (11) is given by Eq. (12) with $E_{n}=E_{m}$ for all $n$ and $m$.
The energy eigenstates of $\hat{H}$ are expressed in the form $|N_{0},N_{S},F,F_{z}\rangle$ for $q=0$, where $N_{S}$ is the number of particles forming singlet pairs, $F$ is the total spin, and $F_{z}$ is the magnetic quantum number. The total number of particles is $N=N_{0}+2N_{S}$. The energy eigenvalue is given by [22]
$E=$ $\varepsilon_{0}N+\frac{c_{0}}{2}N(N-1)$ (14)
$+$ $\frac{c_{1}}{2}[F(F+1)-6N]+\frac{c_{2}}{5}N_{S}(N+N_{0}+3),$
where $c_{0},c_{1}$, and $c_{2}$ are defined as
$c_{0}$ $=\frac{4g_{2}+3g_{4}}{7},$ (15)
$c_{1}$ $=\frac{g_{4}-g_{2}}{7},$
$c_{2}$ $=\frac{7g_{0}-10g_{2}+3g_{4}}{7}.$
We note that the energy eigenstates $|N_{0},N_{S},F,F_{z}\rangle$ are independent of the values of $g_{0},g_{2}$ and $g_{4}$.
Second, let us consider the condition for
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}}=0,$ (16)
where the left-hand side is given by Eq. (10). We will show that such states are mixtures of energy eigenstates with the maximum total spin $F$,
$|F=2N,F_{z}\rangle\equiv|N_{0}=N,N_{S}=0,F=2N,F_{z}\rangle,$ (17)
where $2N\leq F_{z}\leq 2N$. In the following, we write the density matrix in the $N$-particle subspace as $\hat{\rho}_{N}$ and express the density matrix for the total Hilbert space as
$\hat{\rho}=\oplus_{N=0}^{N_{\textrm{ini}}}\hat{\rho}_{N},$ (18)
where $N_{\textrm{ini}}$ is the maximum number of atoms contained in the initial state. We will prove that $\hat{\rho}_{N}$ has the form
$\hat{\rho}_{N}=\sum_{F_{z},F_{z}^{\prime}}c_{F_{z}F_{z}^{\prime}}|F=2N,F_{z}\rangle\langle F=2N,F_{z}^{\prime}|,$ (19)
where $c_{F_{z}F_{z}^{\prime}}$ is a positive semidefinite Hermitian matrix.
To begin with, we prove that Eq. (19) is a sufficient condition for Eq. (16). The states in Eq. (17) are expressed as [22]
$|F=2N,F_{z}\rangle=Z^{-1/2}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle,$ (20)
where $\hat{F}_{-}=2(\hat{a}^{\dagger}_{1}\hat{a}_{2}+\hat{a}_{-2}^{\dagger}\hat{a}_{-1})+\sqrt{6}(\hat{a}^{\dagger}_{0}\hat{a}_{1}+\hat{a}^{\dagger}_{-1}\hat{a}_{0})$, $\Delta F=2N-F_{z}$, $Z=N!(4N)!\Delta F!/(4N-\Delta F)!$ is a normalization constant, and $|vac\rangle$ is the vacuum state. From the form of $\hat{A}_{00}=(\hat{a}_{0}\hat{a}_{0}-2\hat{a}_{1}\hat{a}_{-1}+2\hat{a}_{2}\hat{a}_{-2})/\sqrt{5}$, we find $\hat{A}_{00}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$ and $\hat{A}_{00},\hat{F}_{-}]=0$, which gives $\hat{A}_{00}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$. Similarly, $\hat{A}_{2\mathcal{M}}$ satisfies $\hat{A}_{2\mathcal{M}}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$. From the commutation relations between $\hat{A}_{2\mathcal{M}}$ and $\hat{F}_{-}$,
$\left[\hat{A}_{2,2},\hat{F}_{-}\right]=0,$ (21)
$\left[\hat{A}_{2,1},\hat{F}_{-}\right]=2\hat{A}_{2,2},$
$\left[\hat{A}_{2,0},\hat{F}_{-}\right]=\sqrt{6}\hat{A}_{2,1},$
$\left[\hat{A}_{2,-1},\hat{F}_{-}\right]=2\sqrt{6}\hat{A}_{2,0},$
$\left[\hat{A}_{2,-2},\hat{F}_{-}\right]=4\hat{A}_{2,-1},$
we obtain $\hat{A}_{2\mathcal{M}}(\hat{F}_{-})^{\Delta F}(\hat{a}^{\dagger}_{2})^{N}|vac\rangle=0$. Thus, we have shown that
$\hat{A}_{\mathcal{FM}}|F=2N,F_{z}\rangle=0,$ (22)
for $\mathcal{F}=0$ and 2, and therefore, Eq. (19) satisfies Eq. (16).
Next, we show that Eq. (19) is a necessary condition for Eq. (16). First, we prove the lemma that Eq. (16) leads to
$\hat{A}_{\mathcal{FM}}\hat{\rho}_{N}\hat{A}_{\mathcal{FM}}^{\dagger}=0$ (23)
for $\mathcal{F}=0,2$ and for all possible $N$. We start from the dissipation from the $N=2$ space to the $N=0$ space. From $\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\rho_{0}=\rho_{0}\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}=0$, Eq. (16) gives
$\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}2\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}\hat{A}_{\mathcal{FM}}^{\dagger}=0.$ (24)
Taking the trace of both sides of Eq. (24), we obtain
$\textrm{Tr}\left[\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}2\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}\hat{A}_{\mathcal{FM}}^{\dagger}\right]$ (25)
$=\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}2\textrm{Tr}\left[\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}\right]$
$=0.$
From the positive definiteness of $\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}$, each term of the sum in Eq (LABEL:tracing) must vanish; therefore, Eq. (23) is satisfied for $N=2$. For the dissipation from $N=4$ to $N=2$,
$\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}$ $\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left(2\hat{A}_{\mathcal{FM}}\hat{\rho}_{4}\hat{A}_{\mathcal{FM}}^{\dagger}\right$. (26)
- $\left.\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}_{2}-\hat{\rho}_{2}\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\right)=0$
holds from Eq. (16), where the second and third terms in the bracket vanish according to Eq. (LABEL:tracing). Therefore, by a similar procedure, we can show that Eq. (23) also holds for $N=4$. By mathematical induction, Eq. (23) holds for $N=2,4,...,N_{\textrm{ini}}$. Also, for odd values of $N$, we can similarly show that Eq. (23) holds. Thus, when the density matrix $\hat{\rho}_{N}$ satisfies Eq. (16), Eq. (23) holds for all possible $N$.
In general, the density matrix $\hat{\rho}_{N}$ is written as a mixture of pure states $\{|\psi_{N}^{(k)}\rangle\}_{k=1,2,...,K}$ as
$\hat{\rho}_{N}=\sum_{k=1}^{K}P_{N}^{(k)}|\psi_{N}^{(k)}\rangle\langle\psi_{N}^{(k)}|,$ (27)
where $P_{N}^{(k)}$ is positive. When the density matrix satisfying Eq. (23) is expressed by Eq. (27),
$\hat{A}_{\mathcal{FM}}|\psi_{N}^{(k)}\rangle$ $=0~~~~(\mathcal{F}=0,2),$ (28)
must hold for each pure state $|\psi_{N}^{(k)}\rangle$. The states in Eq. (17) satisfy Eq. (28) as shown in Eq. (22). We can show that no other states satisfy Eq. (28) as follows. Equation (28) leads to
$\sum_{\mathcal{F}=0,2}\frac{g_{\mathcal{F}}}{2V^{\textrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{\mathcal{A}}^{\dagger}_{\mathcal{FM}}\hat{\mathcal{A}}_{\mathcal{FM}}|\psi_{N}^{(k)}\rangle=0.$ (29)
Since the operator on the left-hand side has the same form as the Hamiltonian in Eq. (8) with $\varepsilon_{0}=g_{4}=q=0$, the eigenvalue $\lambda$ is given by
$\lambda=$ $\frac{g_{2}}{7}[2N(2N+1)-F(F+1)]$ (30)
$+$ $\frac{7g_{0}-10g_{2}}{7}N_{S}(N+N_{0}+3)$
from Eq. (14). The quantum numbers $N_{S}$ and $F$ that give $\lambda=0$ are only $N_{S}=0$ and $F=2N$; that is, Eq. (29) is satisfied only by the states in Eq. (17). Therefore, Eq. (28) is satisfied only by the states in Eq. (17). Thus, Eq. (16) is satisfied if and only if the state has the form of Eq. (19). In conclusion, the steady state that satisfies Eqs. (11) and (16) for $q=0$ must have the form of Eq. (18) with Eq. (19).
Finally, we consider the steady state for $q\neq 0$. Since Eq. (16) is independent of the value of $q$, the steady-state density matrix $\hat{\rho}$ must be a mixture of the states given by Eq. (17). However, because of the quadratic Zeeman term in the Hamiltonian,
$\hat{Q}=q\sum_{m=-2}^{2}m^{2}\hat{a}^{\dagger}_{m}\hat{a}_{m},$ (31)
some of the states in Eq. (17) are not the eigenstates of the Hamiltonian; for these states, Eq. (13) does not hold. Among the states in Eq. (17), only the states
$|F=2N,F_{z}$ $=-2N\rangle$ $=\frac{(\hat{a}^{\dagger}_{-2})^{N}}{\sqrt{N!}}|vac\rangle,$ (32)
$|F=2N,F_{z}$ $=2N\rangle$ $=\frac{(\hat{a}^{\dagger}_{2})^{N}}{\sqrt{N!}}|vac\rangle,$
and
$|F=2N,F_{z}$ $=-2N+1\rangle$ $=\frac{\hat{a}^{\dagger}_{-1}(\hat{a}^{\dagger}_{-2})^{N-1}}{\sqrt{(N-1)!}}|vac\rangle,$ (33)
$|F=2N,F_{z}$ $=2N-1\rangle$ $=\frac{\hat{a}^{\dagger}_{1}(\hat{a}^{\dagger}_{2})^{N-1}}{\sqrt{(N-1)!}}|vac\rangle,$
are eigenstates of the Hamiltonian with eigenvalues $E_{q=0}+4qN$ for Eq. (32) and $E_{q=0}+4q(N-1)+q$ for Eq. (33), where $E_{q=0}$ is the eigenvalue for $q=0$. Thus, if the density matrix in Eq. (19) is constructed from the states in Eqs. (32) and (33) such that Eq. (13) is satisfied, we obtain a steady state for $q\neq 0$.
IV NUMERICAL RESULTS
Table 1: Interaction and dissipation parameters $g_{\mathcal{F}}$ and $b_{\mathcal{F}}$ of $^{87}\textrm{Rb}$, taken from Refs. [13] and [20], respectively.
Interaction $10^{-11}\textrm{cm}^{3}/\textrm{s}$ Loss $10^{-11}\textrm{cm}^{3}/\textrm{s}$
$g_{0}/\hbar$ 4.34 $b_{0}$ 0.0099
$g_{2}/\hbar$ 4.59 $b_{2}$ 0.0243
$g_{4}/\hbar$ 5.15 $b_{4}$ -
We study the time evolution of the system by numerically solving the master equation (4) using the fourth-order Runge-Kutta method. The initial state is a Fock state $(\hat{a}^{\dagger}_{0})^{N_{\textrm{ini}}}|\textrm{vac}\rangle/\sqrt{N_{\textrm{ini}}!}$ occupying the $m=0$ state, as in the experiment in Ref. [21]. Here, we restrict ourselves to $N_{\textrm{ini}}=20$, which requires $\sim 10^{9}$ matrix elements to represent the density matrix. We consider an atomic gas of $^{87}\mathrm{Rb}$ confined in an isotropic harmonic potential with a trap frequency $\omega/(2\pi)=2~\mathrm{kHz}$. We use the single-particle ground-state wave function for the single-mode approximation, and the effective volume is given by $V^{\textrm{eff}}\simeq 2.2\times 10^{-13}~\mathrm{cm}^{3}$. The interaction and dissipation parameters are set to the experimentally measured values given in Table 1. For these interaction parameters, the ground state of the Hamiltonian for $q=0$ is the state where all atoms form singlet pairs.
As the system evolves over time, the total density matrix for the system has the form $\hat{\rho}=\oplus_{N=0}^{20}\hat{\rho}_{N}$, where $N$ are even integers. The expectation value of an operator $\hat{A}$ is given by $\langle\hat{A}\rangle=\textrm{Tr}[\hat{A}\hat{\rho}]/\textrm{Tr}[\hat{\rho}$], and that of $\hat{A}$ within the $N$-particle subspace is defined by $\langle\hat{A}\rangle_{N}=\textrm{Tr}[\hat{A}\hat{\rho}_{N}]/\textrm{Tr}[\hat{\rho}_{N}]$. The spin operator in the $j$-direction ($j=x,y,z$) is expressed as $\hat{F}_{j}=\sum_{m,m^{\prime}=-2}^{2}[f_{j}]_{m,m^{\prime}}\hat{a}_{m}^{\dagger}\hat{a}_{m^{\prime}}$, where $f_{j}$ are spin-2 matrices. For the present initial state, $\langle\hat{F}_{j}\rangle$ and $\langle\hat{F}_{j}\rangle_{N}$ always vanish as the system evolves. We define the transverse magnetization as $S_{\perp}=\sqrt{\langle\hat{F}_{x}^{2}+\hat{F}_{y}^{2}\rangle}$. To examine whether the state converges to the fully magnetized steady state in Eq. (18) with Eq. (17), we introduce the normalized spin-squared operator as
$\hat{\textbf{F}}^{2}_{\textrm{norm}}=\frac{\hat{F}_{x}^{2}+\hat{F}_{y}^{2}+\hat{F}_{z}^{2}}{2\hat{N}(2\hat{N}+1)}.$ (34)
To avoid divergence arising from $N=0$, we define the density matrix restricted to the $N>0$ subspace as $\hat{\rho}_{N>0}=\oplus_{N=2}^{20}\hat{\rho}_{N}$. The expectation value of Eq. (34) is taken with respect to this restricted state as
$F^{2}_{\textrm{norm}}=\frac{\textrm{Tr}[\hat{\textbf{F}}^{2}_{\textrm{norm}}\hat{\rho}_{N>0}]}{\textrm{Tr}[\hat{\rho}_{N>0}]}.$ (35)
This quantity becomes unity for the state in Eq. (18) with Eq. (17). Let $P_{N}$ be the probability of finding the system in a state with $N$ particles,
$P_{N}=\frac{\textrm{Tr}[\hat{\rho}_{N}]}{\textrm{Tr}[\hat{\rho}]}$. (36)
We define the probability of measuring a magnetization of $m$ within the $N$-particle subspace as
$Q_{m}^{(N)}=\frac{\textrm{Tr}[\hat{\mathcal{P}}_{m}\hat{\rho}_{N}]}{\textrm{Tr}[\hat{\rho}_{N}]},$ (37)
where $\hat{\mathcal{P}}_{m}$ is an operator describing the projection onto the eigenstates of $\hat{F}_{z}$ with an eigenvalue $m$.
Figure 1: Time evolution for $q=0$. (a) Average particle number $\langle\hat{N}\rangle$ and normalized total magnetization $F^{2}_{\textrm{norm}}$. (b) Particle number distribusions $P_{N}$ at $t=0,0.132$, and $1.76~\mathrm{s}$. (c) Distribusions $Q_{m}^{(N)}$ of the $z$-magnetization at $t=0.132$ and $1.76~\mathrm{s}$. The upper and lower panels show the distributions in the subspaces $N=6$ and $N=16$, respectively. The dotted lines correspond to Eq. (38).
IV.1 Without quadratic Zeeman effect
Figure 1(a) shows the time evolution of the average particle number $\langle\hat{N}\rangle$ and the normalized total magnetization $F^{2}_{\textrm{norm}}$. The average particle number rapidly decreases in the early stages, since the initial state contains an abundance of states with $F<2N$ subject to inelastic decay. After several hundred milliseconds, the states with $F<2N$ are depleted and the decrease in $\langle\hat{N}\rangle$ becomes slow; $\langle\hat{N}\rangle$ finally converges to $\simeq 3$ over long periods. The normalized total magnetization $F^{2}_{\textrm{norm}}$ increases monotonically, as shown in Fig. 1(b), and finally converges to unity, which indicates that the final state is the fully magnetized steady state, given by Eq. (18) with Eq. (17). Because of the rotational symmetry of the system, the expectation values $\langle\hat{F}_{x}\rangle,\langle\hat{F}_{y}\rangle$, and $\langle\hat{F}_{z}\rangle$ of the magnetization remain zero as the system evolves.
Figures 1(b) and 1(c) show the distributions of the number of particles $P_{N}$ and the magnetization $Q_{m}^{(N)}$ at different times. In Fig. 1(b), it can be seen that the distribution becomes broad and its peak shifts toward smaller particle numbers due to inelastic loss. The distribution at $t=1.76~\textrm{s}$ in Fig. 1(c) is similar to the final distribution for the steady state, which has a peak at $N=2$. Figure 1(c) displays the distribution $Q_{m}^{(N)}$ for $N=6$ and $N=16$. We note that $Q_{m}^{(N)}$ is broadened by inelastic loss of particles with $m_{1}+m_{2}\neq 0$, where $m_{1}$ and $m_{2}$ are the magnetic quantum numbers of the lost particles. As a result, a statistical mixture of states with different $m$ is generated.
For comparison, the dotted lines in Fig. 1(c) show
$Q_{m}=\frac{1}{2^{4N}}\binom{4N}{2N+m}$, (38)
which is the distribution of the mean-field states
$|\Psi_{\textrm{MF}}^{(N)}(\phi)\rangle=\hat{R}(\phi)|F=2N,F_{z}=2N\rangle,$ (39)
where $\hat{R}(\phi)=e^{-i\phi\hat{F}_{x}}e^{-i\pi/2\hat{F}_{y}}$ is the rotation operator. These mean-field states correspond to the fully polarized states in the $x$-$y$ directions. The dotted lines for $N=6$ agree well with the results of the numerical simulation, which indicates that the steady state can be approximated by a mixture of $\hat{\rho}_{N}$ as
$\hat{\rho}_{N}=\frac{1}{2\pi}\int d\phi|\Psi_{\textrm{MF}}^{(N)}(\phi)\rangle\langle\Psi_{\textrm{MF}}^{(N)}(\phi)|.$ (40)
This state is a statistical mixture of mean-field states polarized in the $x$-$y$ directions. On the other hand, the distribution $Q_{m}^{(16)}$ in Fig. 1(c) deviates substantially from the dotted line, i.e., the state $\hat{\rho}_{16}$ differs from the mean-field-like state in Eq. (40). However, the probability of observing this state is very small; $P_{16}\sim 10^{-5}$ (see (iii) in Fig. 1(b)).
Figure 2: Time evolution for $q/\hbar=90.6~\textrm{Hz}$ with and without quenching of $q$. In the case with quenching, $q$ is suddenly changed to 0 at $t=14.3~\textrm{ms}$. (a) Transverse magnetization $S_{\perp}$ with quenching (solid line) and without quenching (dotted line). (b) Average particle number $\langle\hat{N}\rangle$ (blue or dark-gray lines) and normalized total magnetization $F^{2}_{\textrm{norm}}$ (red or light-gray lines) with quenching (solid lines) and without quenching (dotted lines). (c) Particle number distribution at $t=1.76~\textrm{s}$ with quench (filled boxes) and without quench (striped boxes). (d) Distribution $Q_{m}^{(N)}$ of $z$-magnetization for $N=6$ and $N=16$ at $t=1.76\textrm{s}$. The dotted lines show Eq. (38).
IV.2 With quadratic Zeeman effect
We apply a quadratic Zeeman term of $q/\hbar=90.6~\textrm{Hz}$ to the system, which corresponds to a magnetic field of $0.178~\textrm{G}$. The dotted line in Fig. 2(a) shows the transverse magnetization $S_{\perp}$. In the early stages of evolution, $S_{\perp}$ rapidly becomes large because of the quadratic Zeeman effect, followed by oscillation. Such behavior has also been observed experimentally [21]. Subsequently, $S_{\perp}$ decays to a small value due to the particle loss. The dotted lines in Fig. 2(b) show the time evolution of the average particle number $\langle\hat{N}\rangle$ and the total magnetization $F^{2}_{\textrm{norm}}$. The average particle number monotonically decreases with time and converges to a small value, since the steady state consists of only a limited set of states, as given by Eqs. (32) and (33). The normalized total magnetization $F^{2}_{\textrm{norm}}$ first oscillates and then increases toward unity. Although the final steady state should have $F^{2}_{\textrm{norm}}=1$, the convergence is very slow, as shown in Fig. 2(b). The striped boxes in Fig. 2(c) show the particle number distribution $P_{N}$ at $t=1.76~\textrm{s}$, which has a peak at $N=2$. This distribution is similar to that for $q=0$ in Fig. 1(b).
Next, we consider the case in which the quadratic Zeeman coefficient is suddenly quenched to $q=0$ at $t=14.3~\textrm{ms}$, at which $S_{\perp}$ reaches its first peak. The solid line in Fig. 2(a) shows that $S_{\perp}$ remains large for a long time. The solid lines in Fig. 2(b) show that $\langle\hat{N}\rangle$ also remains large ($\simeq 13$) and $F^{2}_{\textrm{norm}}$ rapidly converges to unity. The particle number distribution has a peak at $N=16$, as shown by the filled boxes in Fig. 2(c). These behaviors contrast sharply with those in the case without quenching. This is because the fully polarized states in Eq. (17) generated by the quadratic Zeeman term remain entirely in the steady state after $q$ is quenched.
Figure 2(d) shows the distribution $Q_{m}^{(N)}$ of the $z$-magnetization at $t=1.76~\textrm{s}$, which is almost the steady state. For $N=6$, the distribution is in agreement with that for the mean-field-like state in Eq. (40), shown by the dotted line. However, for $N=16$, for which the probability $P_{N}$ is largest, the distribution is nontrivial, exhibiting multiple peaks at $m=0$ and $m=\pm 2$. This characteristic distribution, in which even $m$ is enhanced and odd $m$ is suppressed, suggests a quantum mechanical superposition of fully polarized states oriented in opposite directions ($\theta=-\pi/2$ and $\pi/2$) on the $x$-$y$ plane. This is understood from the fact that the Schödinger-cat-like state given by
$\hat{R}(\pi/2)+\hat{R}(-\pi/2)](\hat{a}_{2}^{\dagger})^{N}|vac\rangle$ (41)
has a distribution $Q_{m}^{(N)}$ that is zero for odd $m$ and is proportional to Eq. (38) for even $m$ [28, 29]. Interestingly, such a nontrivial quantum superposition is achieved despite the system being an open system subject to particle dissipation.
V Conclusions
We have investigated the dynamics of spin-2 bosons with atomic dissipation. In the previous study [21], it was shown experimentally and numerically that the dissipation can enhance the transverse magnetization in the time evolution, and the results of the mean-field calculations were found to agree well with the experiments. By contrast, the present work focused on the quantum many-body nature of a system with a small number of particles. We proved that the steady states for this system are restricted to statistical mixtures of states with maximum total spins, given by Eq. (18) with Eq. (17). We performed numerical simulations using the master equation with and without the quadratic Zeeman term. Without the quadratic Zeeman term, we found that the system indeed reached a steady state with the maximum total spins. When the quadratic Zeeman term was removed in the early stages of time evolution, a relatively large number of particles were found to remain in the steady state. We concluded that this state resembles a Schrödinger-cat-like state, formed by the superposition of mean-field-like states. This suggests that particle dissipation not only destroys the quantum properties but also creates nontrivial quantum many-body states. Although in the present study we considered $^{87}\textrm{Rb}$ atoms, other atomic species, such as $^{23}\textrm{Na}$ atoms, may be experimentally suitable, since their inelastic decay rates are much larger [30, 31, 32] and steady states can be reached much faster.
ACKNOWLEDGMENTS
This work was supported by JSPS KAKENHI Grant Number JP23K03276.